OpenStudy (anonymous):

Simplify sin^2(54)+cos^2(72)

OpenStudy (anonymous):

Please don't just go sin 54 = blah and cos(72) = blah, so blah + blah = blahblah

OpenStudy (anonymous):

@Callisto

OpenStudy (anonymous):

is 54 the actual number or is that in 54 degrees?

OpenStudy (anonymous):

54 degrees

OpenStudy (zzr0ck3r):

So you just want to know how to computte this by hand?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

why did you go into radians?

OpenStudy (anonymous):

so what is x equal to?

OpenStudy (anonymous):

I fluttered that up ignore me.

OpenStudy (anonymous):

Hopefully zrocker can help you X_X

OpenStudy (callisto):

\[sin^2(54)+cos^2(72)\]\[=(cos(90-54))^2+cos^2(72)\]\[=cos^2(36)+cos^2(72)\] That looks nicer

OpenStudy (anonymous):

Oh, wait... let me try now

OpenStudy (callisto):

Okay :|

OpenStudy (anonymous):

cos^2(36) + (2cos^2(36)-1)^2

OpenStudy (anonymous):

am i right?

OpenStudy (callisto):

Yes.

OpenStudy (anonymous):

now... do we evaluate cos(36) or simplify more?

OpenStudy (callisto):

Evaluate.

OpenStudy (anonymous):

ok. is this straightforward or something we fancily work out?

OpenStudy (callisto):

very fancy :|

OpenStudy (callisto):

let x = 36 5x = 180 3x = 180 - 2x sin 3x = sin2x 3 sinx - 4sin^3 x = 2sinx cox Since sinx =/= 0 3- 4 sin^2 x = 2cosx 3- 4 ( 1-cos^2 x) = 2 cos x 4 cos^2 x - 2cosx -1 =0 Now, solve for cosx.

OpenStudy (zzr0ck3r):

wow

OpenStudy (callisto):

@apple_pi Come back :'(

OpenStudy (zzr0ck3r):

So in what class do we learn to play with this trick?

OpenStudy (callisto):

I'm sorry, I don't know. I'm not from the US. But for me, I learnt it two years before university.

OpenStudy (anonymous):

Sorry... I had to go for a while

OpenStudy (anonymous):

Yeah that solved it! Thanks @Callisto

OpenStudy (callisto):

You're welcome. FYI, this post is worth looking! http://openstudy.com/study#/updates/4fe6e718e4b02c91101b52cd

OpenStudy (anonymous):

Got another one: http://openstudy.com/study#/updates/50827553e4b0dab2a5ebddaa